OpenStudy (anonymous):

A projectile of mass m is launched with an initial velocity i making an angle θ with the horizontal as shown below. The projectile moves in the gravitational field of the Earth. Find the angular momentum of the projectile about the origin when the particle is at the following locations. (Use the following as necessary: vi, θ, m, and g for the acceleration due to gravity.) a) at its highest point of trajectory? B) Just before it hits the ground.

OpenStudy (anonymous):

can you draw up the diagram or post it?

OpenStudy (anonymous):

angular momentum = inertia of object X angular velocity and momentum is 'conserved' then we can set up both sides of the 'before event' and 'after event' for an equatoin

OpenStudy (anonymous):

http://www.webassign.net/serpse8/11-p-019.gif

OpenStudy (anonymous):

yikes ^_^ @AllTehMaffs

OpenStudy (anonymous):

Were you able to see the diagram?

OpenStudy (anonymous):

ya, this question is over my head, i messaged somone smatert than me to come and look

OpenStudy (anonymous):

OOOOo okay. This is WEIRD. Okay. one sec.

OpenStudy (anonymous):

I am totally not smarter than you, shush yoself

OpenStudy (anonymous):

thats what I said. but if you can solve the lenghts of a messed up triangle that only has one length given, i know you can handle this, haha

OpenStudy (anonymous):

^_^

OpenStudy (anonymous):

@gkendoyan Do you know the notation \[ \frac{dx}{dt}=\dot x\] ? If no it's not a problem. It just makes things prettier ^_^

OpenStudy (anonymous):

@AllTehMaffs No i don't

OpenStudy (anonymous):

no worries. Okay, so I think this isn't as complicated as it seems. It's just weird. What's the equation for angular momentum?

OpenStudy (anonymous):

L=IW

OpenStudy (anonymous):

yah (also hold on a sec. This is harder than anticipated :P) Have you gotten anywhere with this yet,, or is just one of those "I have no idea" questions

OpenStudy (anonymous):

@AllTehMaffs thats what I was thinking, this problem is weird

OpenStudy (anonymous):

Well i started but couldn't continue.. so basically its one of those i have no idea questions

OpenStudy (anonymous):

where'd ya get to? this gets complicated real quick...

OpenStudy (anonymous):



OpenStudy (anonymous):

hmmm :/

OpenStudy (anonymous):

\[t_{tot}=2\frac{v_0 \sin \theta}{g}\] \[R = v_x t_{tot} = \frac{2v_0^2 \cos \theta \sin \theta}{g}\] \[ h=\frac{(v_0 \sin \theta)^2}{2g}\] \[| \textbf r_{max \ height}|=\sqrt{\Big( \frac{R}{2} \Big)^2+h^2} \\ \ \\ \ \\ \ = \sqrt{\Bigg (\frac{2v_0^2 \cos \theta \sin \theta}{2g}\Bigg)^2+\Bigg( \frac{(v_0 \sin \theta)^2}{2g} \Bigg)^2} \] " t_{tot}=2\frac{v_0 \sin \theta}{g} R = v_x t_{tot} = \frac{2v_0^2 \cos \theta \sin \theta}{g} h=\frac{(v_0 \sin \theta)^2}{2g} | \textbf r_{max \ height}|=\sqrt{\Big( \frac{R}{2} \Big)^2+h^2} \\ \ \\ \ \\ \ = \sqrt{\Bigg (\frac{2v_0^2 \cos \theta \sin \theta}{2g}\Bigg)^2+\Bigg( \frac{(v_0 \sin \theta)^2}{2g} \Bigg)^2} "

OpenStudy (anonymous):

well I guess first, do you see where I got all of those values?

OpenStudy (anonymous):

(I didn't actually write down anything to get to that point. SOrry :P)

OpenStudy (anonymous):

thats fine. I'm trying to follow

OpenStudy (anonymous):

The height is from the fourth equation of motion \[\cancel{v_{fy}^2}^0 = v_{fi}^2+2ah\] and the time is from \[a=\frac{\Delta v_y}{\Delta t}=\frac{\cancel{v_{yf}}^0-v_{yi}}{(1/2) \ t_{tot}}\]

OpenStudy (anonymous):

Angular momnetum is \[L = I \omega\] but also \[\textbf L =\textbf r \times m \textbf v\] Have you studied cross products?

OpenStudy (anonymous):

\[ \textbf r_1 =\frac{R}{2}\hat{ \textbf x } + h \hat{ \textbf y } \\ \ \\ \ \\ \ = \Bigg( \frac{v_0^2 \cos \theta \sin \theta}{g}\Bigg) \hat{ \textbf x } + \Bigg(\frac{(v_0 \sin \theta)^2}{2g} \Bigg) \hat{ \textbf y } \] \[\textbf v = v_{x} \hat {\textbf x} + v_y \hat{ \textbf y}\] Does any of this look familiar? (I'm going at a snail's pace because my brain is slowz)

OpenStudy (anonymous):

Is this looking like it's going in a useful direction @DemolisionWolf ?

OpenStudy (anonymous):

Yeah, I think you need to cross product. Have you done that before?

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Cross_product \[ \textbf L_1 = \textbf r_1 \times m \textbf v_1 = \left|\begin{matrix}\hat{ \textbf x } & \hat{ \textbf y } & \hat{ \textbf k }\\ R/2 & h & 0 \\v_{xi} & 0 & 0 \end{matrix}\right|\] The angular momentum will be about the k direction for both L1 and L2. I gotta run, but I'll be back later. Tell me if this is making sense to you or ^_^

OpenStudy (anonymous):

cross product gives us that at the maximum height, the angular momentum is \[ \textbf L_1 = -hv_{xi} \hat{ \textbf k }\] \[ \textbf L_1 = -\frac{v_0^3 \sin^2 \theta \cos \theta}{2g} \hat{ \textbf k }\] angular momentum is in the z axis pointing into the screen. Makes sense using the right hand rule!

OpenStudy (anonymous):

still need to find L2?